OpenStudy (icalibear):

if the radius is 3x - 2 and the height is x + 3, what is the surface area? use A = 2pi*r^2 + 2pi*rh

OpenStudy (anonymous):

just substitute r = 3x-2 and h =x+3 then simplify

OpenStudy (icalibear):

how do i simplify though

OpenStudy (anonymous):

2pi (3x-2)^2 + 2pi (3x-2)(x+3) expand it first~

OpenStudy (icalibear):

by expand u mean..

OpenStudy (anonymous):

for example (3x-2)^2 = 9x^2 - 12x + 4 (3x-2)(x+3) = 3x^2 +7x-6

OpenStudy (icalibear):

okay and then how do i simplify it

OpenStudy (anonymous):

\[2\pi (3x-2)^2 + 2\pi (3x-2)(x+3)\]\[2\pi (9x^2 - 12x + 4) + 2\pi (3x^2 + 7x -6)\]\[2 \pi (9x^2 + 3x^2 -12x + 7x +4 - 6)\]simplify what's in the brackets then times 2pi to cancel the brackets and u r done

OpenStudy (icalibear):

so the answer is 2pi(12x^2 - 5x - 2)?

OpenStudy (anonymous):

yea or you can times 2pi in if you want \[2 \pi(12x^2 - 5x - 2)= 24 \pi x^2 - 10 \pi x - 4 \pi\]